# Design Optimizer for Soft Growing Robot Manipulators in Three-Dimensional Environments

### v1

@@ 3.5 Constraints
The problem’s constraints are similar to the 2D case presented in the previous work [67], with the respective additions of one more dimension for the rotational angles and one more constraint on obstacle collision that was not previously – and erroneously – considered. Every configuration in the solutions is subject to eight kinematic constraints in addition to the bounds on decision variables (i.e., $\Delta_{\theta}$ and $\Delta_{l}$, which are parameters the user sets prior to manufacturing).
Since the final links (i.e., from the $\epsilon$-th to the $\bar{n}$-th node) are automatically determined to reach the target as described in Section 3.3.1, the orientation angles $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ between the last active link and the target’s orientation segment depends on the kinematic chain produced by the genetic operators. As a result, it may exceed the bounds $\Delta_{\theta}$, as illustrated in Fig. 8(a). We defined four inequality constraints, as shown in Eqn. (13), to mark solutions as infeasible when configurations have $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ beyond these limits. Note that the bound $\Delta_{\theta}$ can also differ between axes, but for the sake of simplicity, we will report it as the same value.
Figure 8: Some examples of infeasible solutions. In (a), excessive steering when reaching the target; in (b), a very small link at the $\epsilon$-th link right after steering that cannot accommodate a gripper device; in (c), not reaching the target with the desired orientation; in (d), an obstacle collision; and in (e) a collision non detected in the configuration but as result of the motion planner, which requires a link to be everted before steering.
@@
Similarly, the length of each configuration is determined by the distance between the $i$-th target and the $\bar{n}_{i}$-th node, which may result in configurations where the length of the final link, $l_{\bar{n}}$, falls below its lower bound limit. In practice, this becomes an issue only when the final link needs to adjust its direction to align with the target’s orientation segment. Considering that these manipulators are likely equipped with a gripping mechanism at the tip (e.g., the tip mount proposed by Jeong et al. [81] or the mechanism to prevent buckling proposed by Coad et al. [82]), it is essential to account for the necessary space to position the gripper properly on the target’s orientation segment for accurate grasping. We formulated the bounds ${\Delta_{l}}^{(L)}$ to account for this constraint. As a result, any solutions featuring configurations where a single link on the target’s orientation segment ($f_{3.1b}=1$, Eqn. (12)) is insufficiently long are considered infeasible, as illustrated in Fig. 8(b). This leads to the definition of a fifth inequality constraint, as expressed in Eqn. (14).
@@
The $\epsilon$-th node, identified as the nearest joint to the target’s orientation segment, is determined using the point-segment distance method (details are reported in Section 1.2 of the supplementary material). Since this process may designate one of the segment’s endpoint as the nearest point to $\epsilon$ – rather than $\epsilon$’s projection on the segment – if that endpoint corresponds to the target, the resulting solution would resemble the configuration shown in Fig. 8(c). Specifically, we refer to the case at line 19 of Algorithm S.3 presented in the supplementary material. To avoid this issue, we established an additional inequality constraint as defined in Eqn. (15). This constraint checks whether the angle width between the end of each configuration $\zeta_{i}$ and the target’s orientation segment $s_{i}$ achieves or exceeds $\pi/2$. Specifically, the target’s orientation segment $s_{i}$ (mentioned in Eqn. (7) as well) has two endpoints: $s_{i,1}$ is the coordinates of its respective target $t_{i,1\rightarrow 3}$, and $s_{i,2}$ is any point on the straight line where $s_{i}$ lies. The angle between these two segments must be smaller than $\pi/2$ for the configuration to be feasible; otherwise, it means that the robot overshot and is trying to align with the target rather than its orientation segment.
@@
Lastly, we impose two constraints related to obstacle collisions. If the environment contains elements that are not targets, they must be treated as obstacles. Any solution that results in a collision with these obstacles is marked as infeasible, as illustrated in Fig. 8(d). In our optimization problem, obstacle collisions are treated as an equality constraint, formulated as shown in Eqn. (16).
@@
The function in Eqn. (16) calculates the number of intersections between each obstacle in the set $o$ and every link of each configuration in $\psi$. The full procedure is reported in Section 1.4 of the supplementary material, and it operates with a complexity of $\mathcal{O}(\overline{n}\cdot|o|\cdot|t|)$. To simplify the process and reduce computational overhead, we model every obstacle as a cylinder due to the ease of detecting intersections between a segment and a cylinder during a collision [90]. More complex obstacles can be modeled by combining multiple cylinders: a pillar is represented by a single cylinder; a wall by a line of cylinders; and a hole in a wall by stacking cylinders in a bottom layer, a top layer, and two taller cylinders at the sides, forming a hole in the middle. During problem setup, users can select a radius slightly larger than that of the obstacle’s circumscribed cylinder to prevent contact between the robot and the obstacle. Note that avoiding obstacles might also be considered as another objective function: one can (i) minimize obstacle collisions or (ii) maximize the distance between any robot link and obstacles. However, the first option would consider a solution with obstacle collision as feasible, which deceives our purposes as we do not want collisions to happen at all. On the other hand, the second option might not work if a target is surrounded by obstacles (e.g., the target is in an open box). Therefore, considering obstacle collisions as a constraint is the best option for our optimization problem.
The last constraint regards a possible obstacle collision when actuating the manipulator – i.e., the static configuration $\zeta_{i}$ does not collide with obstacles, but the actuation to move to $\zeta_{i}$ from any other configuration $\zeta_{j}$ will cause a collision. This is because of how the manipulator is actuated: a link must grow of at least a minimum length (i.e., the lower bound $\Delta^{(L)}_{l}$) before is enabled to steer. Similarly to the constraint expressed in Eqn. (14) and Fig. 8(b), this prevents the robot from steering while the joint is within the tip mount, avoiding possible damage. Fig. 8(e) shows an example: if the configuration requires the robot to grow of a length equal to $\Delta^{(L)}_{l}$ within the obstacle, then we have a collision – note that the figure represents a projection in 2D of a spherical cone. Note that a motion planner could actually find a feasible alternative path without colliding with the obstacle (i.e., by bending one of the previous joints); however, checking this during optimization would require running the motion planner [83], which would require too much in terms of computational resources. Therefore, we simply mark a configuration like this one as infeasible. This constraint is summarized in Eqn. (17), which imposes the number of collisions while steering to be equal to zero. More details are provided in Section 1.5 of the supplementary material.
@@
4 Experimental Results and Discussion
This section illustrates the experiments to (i) assess whether the optimizer can retrieve optimal solutions on tasks with large dimensionality and many obstacles; (ii) determine which EC algorithm is the best fit for solving our design problem among four different options; and (iii) evaluate the robustness of the Rank Partitioning algorithm when included in the loop of different EC algorithms.
@@ 20            $best,f^{best}\leftarrow\textnormal{{rankPartitioning}}(f^{best}\cup F^{pop}_{1})$
4.1 Optimization Methods
We tested the design optimizer on four different EC algorithms: Genetic Algorithm (GA) [91], Particle Swarm Optimization (PSO) [92], Differential Evolution (DE) [93], and Big Bang-Big Crunch algorithm (BBBC) [94]. Each of these EC algorithm has its own parameters, which are described in the following subsections; however, some of the parameters are shared as they all comply with the basic framework of EC algorithms, reported in Algorithm 2: an initial population $pop$ is initialized with $N$ random individuals (line 2), evaluated returning the fitness of each individual $F^{pop}$ penalized in case of constraint violation (line 3), ranked based on the Rank Partitioning algorithm (line 4, runs Algorithm 1), and then it undergoes the generational loop for $G$ iterations (line 5-14) in which variation operators generate a new population of offspring $off$ (line 6), evaluate them returning their fitness $F^{off}$ with penalty in case of constraint violation (line 7), and select which individuals among $pop$ and $off$ will survive and be part of the next generation based on the survival scheme (lines 8-14) – specifically, when elitism is applied, the Rank Partitioning algorithm is executed on both populations (line 9), otherwise only on the offspring (line 12). Note that, with Rank Partitioning, the best individual to be saved and returned by the method is not the individual with minimum/maximum fitness value, but it is the one having rank $=1$ after executing Rank Partitioning between the previous best and the best of the current generation (line 15) – regardless on whether the method is elitist or not. According to preliminary evaluations [^7] For better readability, we do not report any explicit detail of this preliminary study except for the p-values. However, the data can be provided at readers’ request., we set only GA to be elitist (with elitism being significantly better than non-elitist, $p<0.01$), whereas the other EC algorithms were all set to be non-elitist – PSO and DE were found to be significantly better when elitism is not applied ($p<0.01$ and $p=0.0179$, respectively), and no significant difference was observed with BBBC ($p=0.3793$). Table 1 reports the parameters and settings shared among all EC algorithms, with the values we used in the study.
@@ 4.1.2 Particle Swarm Optimization (PSO)
Inertia Weight ($\omega$) $0.6,0.8,1.0$ [101]
Cognitive Constant ($c_{1}$) $0.5,2.5,5.0$ [102, 103]
Social Constant ($c_{2}$) $0.5,2.5,5.0$ [102, 103]
Table 3 reports the parameter settings we used for PSO. The inertia weight $\omega$ is between [0.5, 1.0] after the random $\omega$ factor formula [101]. As for $c_{1}$ and $c_{2}$, Ratnaweera et al. [103] reported the best range for those constants being [0.5, 2.5], with 5.0 being the best parameter suggested by Zhou et al. [102]
@@ 4.1.4 Big Bang-Big Crunch Algorithm (BBBC)
Figure 9: Representation of each task in both isometric (left) and top view (right), along with their respective optimal solutions retrieved by the optimizer.
@@ 4.2 Tasks
We executed each task with a maximum number of nodes $n=20$, bounds on joint rotation set to $\Delta_{\theta}=\pi/4$ on each direction (a limited value emulating hardware constraints), and bounds on link lengths set to $\Delta_{l}\in[25,70]$ to fit the workspace. We repeated each task fifty times for each condition to test all the combinations of parameters discussed in Section 4.1 and summarized in the following Section 4.3.
@@ 4.3 Parameter Combinations
      27 combinations for PSO: 3 values for inertia weight, 3 cognitive constants, and 3 social constants (see Section 4.1.2 and Table 3);
@@ 4.4 Evaluation Metric
Since we are dealing with a five-objective optimization problem (three objectives in which one of them is further divided into two, as described in Section 3.3), we used the rank from Rank Partitioning as the evaluation metric. The Rank Partitioning algorithm, described in Section 3.4, sorts the retrieved solutions based on their objective values according to their predefined priorities; then, it assigns a rank to each of them that linearly increments in ascending order from best to worst – i.e., the best solution gets rank $=1$, the second best gets rank $=2$, and so on.
@@ 4.6 Discussion
4.7 Considerations on Problem Dimensionality
The tasks described in Section 4.2 are optimization problems with high dimensionality in the variable search space. This can be quantified from the fixed value of the maximum link number $n=20$ and the number of targets $|t|$ in the task. Each link accounts for two decision variables per joint (i.e., rotation on $x$ and $y$), which is multiplied for each desired target (i.e., each configuration of the robot aims at one specific target in the environment. The link lengths are shared among all configurations, so they add a constant term. This results in a dimensionality of $2\cdot n\cdot|t|+n$. We can then evaluate the following dimensionalities:
  • 1.
    Task 1 (Fig. 9(a)): $2\cdot 20\cdot 2+20=100$ decision variables;
  • 2.
    Task 2 (Fig. 9(b)): $2\cdot 20\cdot 4+20=180$ decision variables; and
  • 3.
    Task 3 (Fig. 9(c)): $2\cdot 20\cdot 8+20=340$ decision variables.
In the previous 2D work [67], the algorithm was tested only up to 140 decision variables. The tasks analyzed in the current work are inherently more complex due to the third dimension of the environment; yet, as shown by the results reported in Section 4.5, the optimizer can tackle high dimensionality problems while still achieving (sub)optimal designs – especially when the core EC algorithm is GA. It is reasonable to claim that the limit on dimensionality depends on the computational cost of the specific task.
@@ 5 Conclusion
In this work, we presented an extension of our innovative tool to optimize the design of soft growing robots for solving specific tasks. While the previous work was limited only to planar manipulators (i.e., in 2D), the current work completes the tool addressing more realistic 3D scenarios. We used the Rank Partitioning algorithm, introduced in our previous work, to solve a multi-objective optimization problem without compromising among trade-offs. Since this algorithm can theoretically work on every population-based optimizer, we tested it on four different Evolutionary Computation techniques: Genetic Algorithm, Particle Swarm Optimization, Differential Evolution, and Big Bang-Big Crunch algorithm. Results show that the proposed optimizer retains the outstanding performances of its previous 2D version, especially when based on a Genetic Algorithm.

### v2

@@ 3.3 Constraints
The problem’s constraints are similar to the 2D case presented in the previous work [67], with the respective additions of one more dimension for the rotational angles and one more constraint on obstacle collision that was not previously – and erroneously – considered. Every configuration in the solutions is subject to eight kinematic constraints (depicted in Fig. 8) in addition to the decision-variable bounds described in Section 3.1.
Figure 8: Examples of solutions violating constraints – i.e., infeasible. (a) Excessive steering when reaching the target; (b) a small final link that makes it impossible for the robot to carry a gripping mechanism; (c) overshooting the target, resulting in excessive angles; (d) an obstacle collision; and (e) a collision non detected in the configuration but as result of motion, which requires a link to be everted before steering.
The first four constraints concern the angles at the $\epsilon$-th node, which are set by the evaluation function such that, from that point on, the robot grows in the direction of the target – as described in Section 3.2. Since these angles are geometrically calculated, there is no guarantee that they respect the bounds $\Delta_{\theta}$, as illustrated in Fig. 8(a); therefore, we included the four inequality constraints defined in Eqn. (11). Note that the bound $\Delta_{\theta}$ can also differ between axes, but for the sake of simplicity, we will report it as the same value.
@@
The fifth constraint considers the case in which the length of the last link is smaller than the lower bound ${\Delta_{l}}^{(L)}$, as shown in Fig. 8(b). This is an issue only when the final link needs to adjust its direction to align with the target’s orientation segment – i.e., there is only one link in on the segment, or $f_{3.1b}=1$, Eqn. (7): these soft growing robots are likely equipped with a gripper device that moves with the end effector along the robot’s body (e.g., the tip mount proposed by Jeong et al. [81] or the mechanism to prevent buckling proposed by Coad et al. [82]). For this reason, it is essential to account for the necessary space to position this gripper properly on the target’s orientation segment for accurate grasping. The respective inequality constraint is expressed in Eqn. (12).
@@
The sixth constraint handles an edge case of the point-segment distance method of Eqn. (5) – details of the procedure are reported in Section 1.2 of the Supplementary Material. Since this process may designate one of the segment’s endpoint as the nearest point to $\epsilon$ – rather than $\epsilon$’s projection on the segment – if that endpoint corresponds to the target, the resulting solution would resemble the configuration shown in Fig. 8(c). Specifically, we refer to the case at line 19 of Algorithm S.3 presented in the Supplementary Material. To avoid this issue, we established an additional inequality constraint as defined in Eqn. (13). This constraint checks whether the angle width between the end of each configuration $\zeta_{i}$ and the target’s orientation segment $s_{i}$ achieves or exceeds $\pi/2$. Specifically, the target’s orientation segment $s_{i}$ (mentioned in Eqn. (5) as well) has two endpoints: $s_{i,1}$ is the coordinates of its respective target $t_{i,1\rightarrow 3}$, and $s_{i,2}$ is any point on the straight line where $s_{i}$ lies. The angle between these two segments must be smaller than $\pi/2$ for the configuration to be feasible; otherwise, it means that the robot overshot and is trying to align with the target rather than its orientation segment.
@@
Lastly, we impose two constraints related to obstacle collisions. The first one imposes that no collision is generated in the solution, as a robot colliding with an obstacle would be an infeasible solution – as illustrated in Fig. 8(d). The equality constraint in Eqn. (14) evaluates the intersections between the links in $\psi$ and the obstacles in the set $o$. The full procedure is reported in Section 1.4 of the Supplementary Material, and it operates with a complexity of $\mathcal{O}(\overline{n}\cdot|o|\cdot|t|)$. To simplify the process and reduce computational overhead, we model every obstacle as a cylinder due to the ease of detecting intersections between a segment and a cylinder during a collision [90]. More complex obstacles can be modeled by combining multiple cylinders: a pillar is represented by a single cylinder; a wall by a line of cylinders; and a hole in a wall by stacking cylinders in a bottom layer, a top layer, and two taller cylinders at the sides, forming a hole in the middle. During problem setup, users can select a larger radius than that of the obstacle’s circumscribed cylinder to prevent contact between the obstacle and the manipulator. Note that avoiding obstacles might also be considered as another objective function: one can (i) minimize obstacle collisions or (ii) maximize the distance between any robot link and obstacles. However, the first option would consider a solution with obstacle collision as feasible, which deceives our purposes as we do not want collisions to happen at all. On the other hand, the second option might not work if a target is surrounded by obstacles (e.g., the target is in an open box). Therefore, considering obstacle collisions as a constraint is the best option for our optimization problem.
@@
The last constraint regards a possible obstacle collision when actuating the manipulator – i.e., the static configuration $\zeta_{i}$ does not collide with obstacles, but the actuation to move to $\zeta_{i}$ from any other configuration $\zeta_{j}$ will cause a collision. This is because of how the manipulator is actuated: a link must grow of at least a minimum length (i.e., the lower bound $\Delta^{(L)}_{l}$) before is enabled to steer. Similarly to the constraint expressed in Eqn. (12) and Fig. 8(b), this prevents the robot from steering while the joint is within the tip mount, avoiding possible damage. Fig. 8(e) shows an example: if the configuration requires the robot to grow of a length equal to $\Delta^{(L)}_{l}$ within the obstacle, then we have a collision – note that the figure represents a projection in 2D of a spherical cone. A motion planner could actually find a feasible alternative path without colliding with the obstacle (i.e., by bending one of the previous joints); however, checking this during optimization would require running the motion planner [83], which would require too much in terms of computational resources. Therefore, we simply mark a configuration like this one as infeasible. This constraint is summarized in Eqn. (15), which imposes the number of collisions while steering to be equal to zero. More details are provided in Section 1.5 of the Supplementary Material.
@@
4 Experiments
We ran a set of experiments to (i) assess whether the optimizer can retrieve optimal solutions on tasks with large dimensionality and many obstacles; (ii) determine which EC algorithm is the best fit for solving our design problem among four different options; and (iii) evaluate the robustness of the Rank Partitioning algorithm when included in the loop of different EC algorithms.
4.1 Optimization Methods
To solve this design optimization problem, we employ the Rank Partitioning method, which was introduced in the optimizer for the 2D case [67]. Rank partitioning is a non-parametric survival strategy for multi-objective EC algorithms (or generically for population-based optimization methods) that ranks each individual based on a predefined order of the objectives. It sorts the population for the first objective (i.e., the one with the highest priority) and then creates partitions based on their objective rank; then, it sorts each partition for the second objective, and so on. Finally, it returns only one solution: the optimal based on the given order of priorities – specifically, the priority order we selected for our objectives is $f_{1-2}$ (to ensure the robot reaches the targets), $f_{3.1a}$ (to avoid configurations roaming before reaching the correct orientation), $f_{3.2}$ (to prevent wavy configurations), $f_{3.1b}$ and $f_{3.2}$ (to minimize material waste). The advantages of such a method are: (i) it allows us to retrieve the optimal solution based on our preference without compromising trade-offs; (ii) there is no need to retrieve the full Pareto optimal front, saving computational resources; and (iii) it outperforms weighted sum based methods while without featuring their disadvantages [78, 89]. In this section, we seek evidence for a fourth advantage: this method can be applied to any EC algorithm, or any multi or many-objective optimization (i.e., with more than three objectives [86]). More details about Rank Partitioning are reported in Section 2 of the Supplementary Material or in our previous work [67].
@@ 20            $best,f^{best}\leftarrow\textnormal{{rankPartitioning}}(f^{best}\cup F^{pop}_{1})$
We tested the design optimizer on four different EC algorithms: Genetic Algorithm (GA) [91], Particle Swarm Optimization (PSO) [92], Differential Evolution (DE) [93], and Big Bang-Big Crunch algorithm (BBBC) [94]. Each of these EC algorithm has its own parameters, which are described in the following subsections; however, some of the parameters are shared as they all comply with the basic framework of EC algorithms, reported in Algorithm 1: an initial population $pop$ is initialized with $N$ random individuals (line 2), evaluated returning the fitness of each individual $F^{pop}$ penalized in case of constraint violation (line 3), ranked based on the Rank Partitioning algorithm (line 4, runs Algorithm S.11 of the Supplementary Material), and then it undergoes the generational loop for $G$ iterations (line 5-14) in which variation operators generate a new population of offspring $off$ (line 6), evaluate them returning their fitness $F^{off}$ with penalty in case of constraint violation (line 7), and select which individuals among $pop$ and $off$ will survive and be part of the next generation based on the survival scheme (lines 8-14) – specifically, when elitism is applied, the Rank Partitioning algorithm is executed on both populations (line 9), otherwise only on the offspring (line 12). Note that, with Rank Partitioning, the best individual to be saved and returned by the method is not the individual with minimum/maximum fitness value, but it is the one having rank $=1$ after executing Rank Partitioning between the previous best and the best of the current generation (line 15) – regardless on whether the method is elitist or not. According to preliminary evaluations [^2] For better readability, we do not report any explicit detail of this preliminary study except for the p-values. However, the data can be provided at readers’ request., we set only GA to be elitist (with elitism being significantly better than non-elitist, $p<0.01$), whereas the other EC algorithms were all set to be non-elitist – PSO and DE were found to be significantly better when elitism is not applied ($p<0.01$ and $p=0.0179$, respectively), and no significant difference was observed with BBBC ($p=0.3793$). Table 4 reports the parameters and settings shared among all EC algorithms, with the values we used in the study.
@@ 4.1.2 Particle Swarm Optimization (PSO)
Population Size ($N$) $500$
Max Number of Generations ($G$) $500$
Constraint Handling Static Penalty [95, 84] $R=100$
Rank Partitioning Bin Size for $f_{1-2}=0.5$ for $f_{3.2}=5.0$
Table LABEL:tab:param_pso reports the parameter settings we used for PSO. The inertia weight $\omega$ is between [0.5, 1.0] after the random $\omega$ factor formula [101]. As for $c_{1}$ and $c_{2}$, Ratnaweera et al. [103] reported the best range for those constants being [0.5, 2.5], with 5.0 being the best parameter suggested by Zhou et al. [102]
@@ 4.1.4 Big Bang-Big Crunch Algorithm (BBBC)
Figure 9: Representation of each task in both isometric (left) and top view (right), along with their respective solutions.
@@ 4.2 Tasks
We executed each task with a maximum link number $n=20$, joint rotation bounds set to $\Delta_{\theta}=\pi/4$ on each direction (a limited value emulating the physical robot’s capabilities), and link length bounds set to $\Delta_{l}\in[25,70]$ to fit the workspace. We repeated each task fifty times.
Concerning dimensionality, each rotational joint comprises two decision variables (i.e., rotation on $x$ and $y$) multiplied for each $|t|$ target (i.e., each configuration of the robot aims at one specific target in the environment). Configurations share link lengths, so they add a fixed constant value $n$. The overall dimensionality is $2\cdot|t|\cdot n+n$, which is upper-bounded by $2\cdot|t|\cdot n^{\Delta_{\theta}}+n^{\Delta_{l}}$. Specifically, the tasks we tested have the following dimensionalities:
  • 1.
    Task 1: $2\cdot 20\cdot 2+20=100$ decision variables;
  • 2.
    Task 2: $2\cdot 20\cdot 4+20=180$ decision variables; and
  • 3.
    Task 3: $2\cdot 20\cdot 8+20=340$ decision variables.
@@ 4.3 Parameter Combinations
      27 combinations for PSO: 3 values for inertia weight, 3 cognitive constants, and 3 social constants (see Section 4.1.2 and Table LABEL:tab:param_pso);
@@ 4.4 Evaluation Metric
Since we are dealing with a five-objective optimization problem (three objectives in which one of them is further divided into two, as described in Section 3.2), we used the rank from Rank Partitioning as the evaluation metric. The Rank Partitioning algorithm, described in Section 2 of the Supplementary Material, sorts the retrieved solutions based on their objective values according to their predefined priorities; then, it assigns a rank to each of them that linearly increments in ascending order from best to worst – i.e., the best solution gets rank $=1$, the second best gets rank $=2$, and so on.
@@ 4.6 Discussion
Lastly, we discuss the potential of the optimizer to solve problems with high dimensionality – with reference to the values reported in Section 4.2. In the previous 2D work [67], the algorithm was tested only up to 140 decision variables. The tasks analyzed in the current work are inherently more complex due to the third dimension of the environment; yet, as shown by the results reported in Section 4.5, the optimizer can tackle high dimensionality problems while still achieving (sub)optimal designs – especially when the core EC algorithm is GA.
@@ 5 Conclusion
This work presents an extension of our tool for soft-growing design optimization [67], offered to robotic enthusiasts when facing a specific task. While the previous work was limited only to planar manipulators (i.e., in 2D), the current work completes the tool addressing more realistic 3D scenarios. We used the Rank Partitioning algorithm, introduced in our previous work, to solve a multi-objective optimization problem without compromising among trade-offs. Since this algorithm can theoretically work on every population-based optimizer, we tested it on four different Evolutionary Computation techniques: Genetic Algorithm, Particle Swarm Optimization, Differential Evolution, and Big Bang-Big Crunch algorithm. Results show that the proposed optimizer retains the outstanding performances of its previous 2D version, especially when based on a Genetic Algorithm.